# Joint Beamforming and Offloading Design for Integrated Sensing, Communication and Computation System

### 04 Jan 2024 v1

@@ II System Model
We consider a three-tier ISCC system, as depicted in Fig. 1, comprising a cloud server, $M$ base stations (BSs) each equipped with $N_{t}$ transmit antennas and $N_{r}$ receive antennas, and $L$ UTs equipped with $K$ antennas. Here we assume $N_{t}=N_{r}=N$ and all UTs operate on the same frequency band. The BSs are assumed to be able to provide edge computing capabilities for UTs and optionally forward the sensing tasks to cloud through optical fibers. The UTs could be UAVs, vehicles, or robots engaged in diverse sensing tasks such as target detection/estimation and collision avoidance. By employing the ISAC technique, UTs can simultaneously execute sensing and task offloading. In particular, we consider the ISAC signals are transmitted via the shared antenna array structure in [2]. The received signal at BS $m$ can be expressed as
$\mathbf{y}_{m}=\mathbf{H}_{m,i}^{H}\mathbf{W}_{i}\mathbf{s}_{i}+\sum_{l=1,l\neq i}^{L}\mathbf{H}_{m,l}^{H}\mathbf{W}_{l}\mathbf{s}_{l}+\mathbf{z}_{m}$ (1)
where $\mathbf{s}_{i}\in\mathbb{C}^{d}$ denotes $d$ parallel data streams from UT $i$, $\mathbf{W}_{i}\in\mathbb{C}^{K\times d}$ denotes the transmit beamforming matrix of UT $i$, $\mathbf{H}_{m,i}=\in\mathbb{C}^{K\times N}$ denotes the rayleigh channel matrix from UT $i$ to BS $m$, and $\mathbf{z}_{m}\sim\mathcal{CN}(0,{\sigma_{c}}^{2}\mathbf{I}_{N})$ denotes the comlex Gaussian noise at BS $m$. The data rate from UT $i$ to BS $m$ is given by
@@
$\mathbf{r}_{i}=\alpha_{i}\mathbf{a}^{*}({\theta}_{i})\mathbf{a}^{H}({\theta}_{i})\mathbf{W}_{i}\mathbf{s}_{i}+\sum_{l=1,l\neq i}^{L}\widehat{\mathbf{H}}_{l,i}\mathbf{W}_{l}\mathbf{s}_{l}+\hat{\mathbf{z}}_{i},$ (4)
where $\widehat{\mathbf{H}}_{l,i}$ represents the interference channel from UT $l$ to UT $i$, $\hat{\mathbf{z}}_{i}\sim\mathcal{CN}(0,{\sigma_{r}}^{2}\mathbf{I}_{K})$ is the noise vector at UT $i$, and $\alpha_{i}=\sqrt{{\rho}/{{d^{r}_{i}}^{2}}\times{\zeta}_{i}/{{d^{r}_{i}}^{2}}}$ with $\rho$ being the path loss at the reference distance of $d_{0}=1$m, ${\zeta}_{i}$ being the radar cross-section (RCS) of UT $i$’s sensing target, and $d^{r}_{i}$ being the distance from UT $l$ to the sensing target. $\mathbf{a}({\theta}_{i})=[1,{\rm e}^{j2\pi\Delta\sin(\theta_{i})/\lambda},\cdots,{\rm e}^{j2\pi(N-1)\Delta\sin(\theta_{i})/\lambda}]^{T}\in\mathbb{C}^{K\times 1}$ is the steering vector with $\Delta$ being the antenna spacing, ${\theta}_{i}$ being the angle of arrival corresponding to UT $i$’s target, and $\lambda$ being the wavelength. By defining $\mathbf{A}({\theta}_{i})=\mathbf{a}^{*}({\theta}_{i})\mathbf{a}^{H}({\theta}_{i})$, UT $i$’s echo SINR can be expressed as
@@ IV-A Beamforming Matrix Optimization
From (2), we observe that the optimization variables $\mathbf{W}_{i}$ within $R_{mi}$ are located inside the function $\log\det(\cdot)$, rendering it non-convex. We begin by transforming $R_{mi}$ into a concave form. By denoting the $j$-th column of the beamforming matrix $\mathbf{W}_{i}$ as $\mathbf{w}^{j}_{i}$, $R_{m,i}$ can be expressed as the summation of different data streams received by the MEC server [7], i.e.,
$\widehat{R}_{m,i}=\sum_{j=1}^{d}B\log(1+\mathbf{w}^{jH}_{i}\mathbf{H}_{m,i}^{H}\mathbf{N}^{-1}_{m,i}\mathbf{H}_{m,i}\mathbf{w}^{j}_{i}).$ (17)
By applying the quadratic transform [11] to each data stream, (17) can be futher reformulated as
$\tilde{R}_{m,i}=\max_{\bm{z}_{ji}}\sum_{j=1}^{d}B\log(1+2\Re\{\bm{z}^{H}_{j,i}\mathbf{H}_{m,i}\mathbf{w}^{j}_{i}\}-\bm{z}^{H}_{j,i}\mathbf{N}_{m,i}\bm{z}_{j,i}),$ (18)
where $\bm{z}_{j,i}$ are auxiliary variables. Note that the SINR term in (17) has been converted into a concave function with respect to $\mathbf{w}^{j}_{k}$ in (18). Therefore, $\bm{z}_{j,i}$’s and $\mathbf{w}_{i}$’s can be optimized in an iterative way. In each iteration. $\bm{z}_{j,i}$ is optimally updated by
$\bm{z}_{j,i}=\mathbf{N}^{-1}_{m,i}\mathbf{H}_{m,i}\mathbf{w}^{j}_{i},$ (19)
@@
which can be solved via the interior point method [12]. By iteratively handling problem (22) and updating $\bm{z}_{j,i}$ as (19), the original sum of ratios problem (15) can be solved.
@@
2:Initialize offloading decision variables $\bm{a}_{m},\bm{b}_{m}$, auxiliary variables $\bm{z}_{j,i}$ and transmit beamforming matrices $\widetilde{\mathbf{W}}_{i}$ to a feasible value.
@@
5:         Update $\mathbf{w}_{i}^{j}$’s by solving problem (22).
6:         Update $\bm{z}_{j,i}$ by equation (19).
@@ IV-C Overall Algorithm
$\tilde{f}^{(iter+1)}(\mathbf{w}^{j(iter+1)},\bm{z}^{(iter)})\leq\tilde{f}^{(iter)}(\mathbf{w}^{j((iter))},\bm{z}^{(iter)}),$ (24)
@@
For the update of $\bm{z}^{(iter)}_{j,i}$, we first re-express $\tilde{R}_{m,i}$ as $B\log(1+\bm{z}^{H}_{j,i}\mathbf{H}_{m,i}\mathbf{w}^{j}_{i}+\mathbf{w}^{jH}_{i}\mathbf{H}^{H}_{m,i}\bm{z}_{j,i}-\bm{z}^{H}_{m,i}\mathbf{N}_{m,i}\bm{z}_{j,i})$, which is further rewritten as $B\log(1+\mathbf{w}^{jH}_{i}\mathbf{H}_{m,i}^{H}\mathbf{N}^{-1}_{m,i}\mathbf{H}_{m,i}\mathbf{w}^{j}_{i}-(\bm{z}^{H}_{j,i}-\mathbf{w}^{jH}_{i}\mathbf{H}^{H}_{m,i}\mathbf{N}^{-1}_{m,i})\mathbf{N}_{m,i}(\bm{z}_{j,i}-\mathbf{N}^{-1}_{m,i}\mathbf{H}_{m,i}\mathbf{w}^{j}_{i})$. It is easy to observe that by updating $\bm{z}^{(iter)}_{j,i}$ via (19), the left side of constraint (22a) is non-decreasing. Thus, we have
$\small\tilde{R}^{(iter+1)}_{j,i}(\mathbf{w}^{j(iter+1)}_{i},\bm{z}^{(iter+1)}_{j,i})\geq\tilde{R}^{(iter+1)}_{j,i}(\mathbf{w}^{j(iter+1)}_{i},\bm{z}^{(iter)}_{j,i}).$ (25)
@@
$\small\tilde{f}^{(iter+1)}(\mathbf{w}^{j(iter+1)}_{,}\bm{z}^{(iter+1)})\leq\tilde{f}^{(iter)}(\mathbf{w}^{j(iter+1)},\bm{z}^{(iter+1)}).$ (26)
@@
The computational complexity of Algorithm 1 comprises two main components involving solving subproblems (22) and (23). Refering to [14], the overall complexity of Algorithm 1 is $\mathcal{O}\left(N_{iter}(K^{3}L^{3}d^{3}+M^{3}L^{3})\right)$, where $N_{iter}$ represents the number of iterations.
@@ V Simulation Results
In this section, we demonstrate the effectiveness of the proposed three-tier computation offloading scheme integrating beamforming optimization (THCO-BO). We set the antennas numbers of UTs and BSs as $K=12$ and $N=16$, respectively. The computing frequency of UTs, MEC server, and cloud server are set as $f_{i}^{L}=0.1$ Gcycles/s, $f_{i}^{M}=3$ Gcycles/s, and $f_{i}^{C}=10$ Gcycles/s, respectively. $\beta$ is set as 40 cycles/bit [10]. The numbers of BSs and UTs are $M=3$ and $L=9$, respectively. Three other schemes, including two-tier (MEC and UTs) computation offloading scheme with beamforming optimization (TTCO-BO) [10], three-tier computation offloading scheme with maximal ratio communication (THCO-MRC, $\mathbf{W}_{i}=\sqrt{P_{c}}\frac{\mathbf{H}_{mi}[1:d]}{||\mathbf{H}_{mi}||}$) and three-tier computation offloading scheme with maximal ratio sensing (THCO-MRS, $\mathbf{W}_{i}=\sqrt{P_{c}}\frac{\mathbf{A}^{H}({\theta}_{i})[1:d]}{||\mathbf{A}({\theta}_{i})||}$) are carried out for comparisons.

### 26 Jan 2024 v2

@@ II System Model
We consider a three-tier ISCC system, as depicted in Fig. 1, comprising a cloud server, $M$ base stations (BSs) each equipped with $N$ transmit antennas and $N_{r}$ receive antennas, and $L$ UTs equipped with $K$ antennas. Here we assume $N_{t}=N_{r}=N$ and all UTs operate on the same frequency band. The BSs are assumed to be able to provide edge computing capabilities for UTs and optionally forward the sensing tasks to cloud through optical fibers. The UTs could be UAVs, vehicles, or robots engaged in diverse sensing tasks such as target detection/estimation and collision avoidance. By employing the ISAC technique, UTs can simultaneously execute sensing and task offloading. In particular, we consider the ISAC signals are transmitted via the shared antenna array structure in [2]. The received signal at BS $m$ can be expressed as
$\mathbf{y}_{m}=\mathbf{H}_{m,i}^{H}\mathbf{w}_{i}s_{i}+\sum_{l=1,l\neq i}^{L}\mathbf{H}_{m,l}^{H}\mathbf{w}_{l}s_{l}+\mathbf{z}_{m}$ (1)
where $s_{i}$ denotes the data stream from UT $i$, $\mathbf{w}_{i}\in\mathbb{C}^{K\times d}$ denotes the transmit beamforming matrix of UT $i$, $\mathbf{H}_{m,i}=\in\mathbb{C}^{K\times N}$ denotes the rayleigh channel matrix from UT $i$ to BS $m$, and $\mathbf{z}_{m}\sim\mathcal{CN}(0,{\sigma_{c}}^{2}\mathbf{I}_{N})$ denotes the comlex Gaussian noise at BS $m$. The data rate from UT $i$ to BS $m$ is given by
@@
$\mathbf{r}_{i}=\alpha_{i}\mathbf{a}^{*}({\theta}_{i})\mathbf{a}^{H}({\theta}_{i})\mathbf{w}_{i}s_{i}+\sum_{l=1,l\neq i}^{L}\widehat{\mathbf{H}}_{l,i}\mathbf{w}_{l}s_{l}+\hat{\mathbf{z}}_{i},$ (4)
where $\widehat{\mathbf{H}}_{l,i}$ represents the interference channel from UT $l$ to UT $i$, $\hat{\mathbf{z}}_{i}\sim\mathcal{CN}(0,{\sigma_{r}}^{2}\mathbf{I}_{K})$ is the noise vector at UT $i$, and $\alpha_{i}=\sqrt{{\rho}/{{d^{r}_{i}}^{2}}\times{\zeta}_{i}/{{d^{r}_{i}}^{2}}}$ with $\rho$ being the path loss at the reference distance of $d_{0}=1$m, ${\zeta}_{i}$ being the radar cross-section (RCS) of UT $i$’s sensing target, and $d^{r}_{i}$ being the distance from UT $l$ to the sensing target. $\mathbf{a}({\theta}_{i})=[1,{\rm e}^{2\pi\Delta\sin(\theta_{i})/\lambda},\cdots,{\rm e}^{2\pi(N-1)\Delta\sin(\theta_{i})/\lambda}]^{T}\in\mathbb{C}^{K\times 1}$ is the steering vector with $\Delta$ being the antenna spacing, ${\theta}_{i}$ being the angle of arrival corresponding to UT $i$’s target, and $\lambda$ being the wavelength. By defining $\mathbf{A}({\theta}_{i})=\mathbf{a}^{*}({\theta}_{i})\mathbf{a}^{H}({\theta}_{i})$, UT $i$’s echo SINR can be expressed as
@@ IV-A Beamforming Matrix Optimization
From (2), we observe that the optimization variables $\mathbf{w}_{i}$ within $R_{mi}$ are located inside the function $\log\det(\cdot)$, rendering it non-convex. We begin by transforming $R_{mi}$ into a concave form. $R_{m,i}$ can be expressed as [7]
$\widehat{R}_{m,i}=B\log(1+\mathbf{w}^{H}_{i}\mathbf{H}_{m,i}\mathbf{N}^{-1}_{m,i}\mathbf{H}_{m,i}^{H}\mathbf{w}_{i}).$ (17)
By applying the quadratic transform [11], (17) can be futher reformulated as
$\tilde{R}_{m,i}=\max_{\bm{z}_{i}}B\log(1+2\Re\{\bm{z}^{H}_{i}\mathbf{H}^{H}_{m,i}\mathbf{w}_{i}\}-\bm{z}^{H}_{i}\mathbf{N}_{m,i}\bm{z}_{i}),$ (18)
where $\bm{z}_{i}$ are auxiliary variables. Note that the SINR term in (17) has been converted into a concave function with respect to $\mathbf{w}_{k}$ in (18). Therefore, $\bm{z}_{i}$’s and $\mathbf{w}_{i}$’s can be optimized in an iterative way. In each iteration. $\bm{z}_{i}$ is optimally updated by
$\bm{z}_{i}=\mathbf{N}^{-1}_{m,i}\mathbf{H}_{m,i}\mathbf{w}_{i},$ (19)
@@
which can be solved via the interior point method [12]. By iteratively handling problem (22) and updating $\bm{z}_{i}$ as (19), the original sum of ratios problem (15) can be solved.
@@
2:Initialize offloading decision variables $\bm{a}_{m},\bm{b}_{m}$, auxiliary variables $\bm{z}_{i}$ and transmit beamforming matrices $\widetilde{\mathbf{w}}_{i}$ to a feasible value.
@@
5:         Update $\mathbf{w}_{i}$’s by solving problem (22).
6:         Update $\bm{z}_{i}$ by equation (19).
@@ IV-C Overall Algorithm
$\tilde{f}^{(iter+1)}(\mathbf{w}^{(iter+1)},\bm{z}^{(iter)})\leq\tilde{f}^{(iter)}(\mathbf{w}^{((iter))},\bm{z}^{(iter)}),$ (24)
@@
For the update of $\bm{z}^{(iter)}_{i}$, we first re-express $\tilde{R}_{m,i}$ as $B\log(1+\bm{z}^{H}_{i}\mathbf{H}_{m,i}\mathbf{w}_{i}+\mathbf{w}^{H}_{i}\mathbf{H}^{H}_{m,i}\bm{z}_{i}-\bm{z}^{H}_{m,i}\mathbf{N}_{m,i}\bm{z}_{i})$, which is further rewritten as $B\log(1+\mathbf{w}^{H}_{i}\mathbf{H}_{m,i}^{H}\mathbf{N}^{-1}_{m,i}\mathbf{H}_{m,i}\mathbf{w}_{i}-(\bm{z}^{H}_{i}-\mathbf{w}^{H}_{i}\mathbf{H}^{H}_{m,i}\mathbf{N}^{-1}_{m,i})\mathbf{N}_{m,i}(\bm{z}_{i}-\mathbf{N}^{-1}_{m,i}\mathbf{H}_{m,i}\mathbf{w}_{i})$. It is easy to observe that by updating $\bm{z}^{(iter)}_{i}$ via (19), the left side of constraint (22a) is non-decreasing. Thus, we have
$\small\tilde{R}^{(iter+1)}_{i}(\mathbf{w}^{(iter+1)}_{i},\bm{z}^{(iter+1)}_{i})\geq\tilde{R}^{(iter+1)}_{i}(\mathbf{w}^{(iter+1)}_{i},\bm{z}^{(iter)}_{i}).$ (25)
@@
$\small\tilde{f}^{(iter+1)}(\mathbf{w}^{(iter+1)}_{,}\bm{z}^{(iter+1)})\leq\tilde{f}^{(iter)}(\mathbf{w}^{(iter+1)},\bm{z}^{(iter+1)}).$ (26)
@@
The computational complexity of Algorithm 1 comprises three main components involving update $\bm{z}_{i}$ and solving subproblems (22) and (23). Refering to [14], the overall complexity of Algorithm 1 is $\mathcal{O}\left(N_{iter}(K^{3}L^{3}+M^{3}L^{3}+N^{3}L+KN^{2}L)\right)$, where $N_{iter}$ represents the number of iterations.
@@ V Simulation Results
In this section, we demonstrate the effectiveness of the proposed three-tier computation offloading scheme integrating beamforming optimization (THCO-BO). We set the antennas numbers of UTs and BSs as $K=12$ and $N=16$, respectively. The computing frequency of UTs, MEC server, and cloud server are set as $f_{i}^{L}=0.1$ Gcycles/s, $f_{i}^{M}=3$ Gcycles/s, and $f_{i}^{C}=10$ Gcycles/s, respectively. $\beta$ is set as 40 cycles/bit [10]. The numbers of BSs and UTs are $M=3$ and $L=9$, respectively. Three other schemes, including two-tier (MEC and UTs) computation offloading scheme with beamforming optimization (TTCO-BO) [10], three-tier computation offloading scheme with maximal ratio communication (THCO-MRC, $\mathbf{w}_{i}=\sqrt{P_{c}}\frac{\mathbf{H}_{mi}[\lambda_{max}]}{||\mathbf{H}_{mi}||}$) and three-tier computation offloading scheme with maximal ratio sensing (THCO-MRS, $\mathbf{w}_{i}=\sqrt{P_{c}}\frac{\mathbf{a}({\theta}_{i})}{||\mathbf{a}({\theta}_{i})||}$) are carried out for comparisons, where $\lambda_{max}$ represents the column corresponding to the largest eigenvalue of $\mathbf{H}_{mi}$.